what is
Trains A and B are traveling in the same direction on parallel tracks. train a is traveling at 100 miles per hour and train b is traveling at 125 miles per hour. Train A passes a station at 2:10 A.M If train B passes the same station at 2:22 am, at what time will train b catch up to train A?
When will train B catch up with train A?

Answer by checkley77(12569)

100t=125(t-12/60)
100t=125t-.2*125
100t-125t=-25
25t=25
t=25/25
t=1 hour after train leave they will meet.
Proof:
100*1=125(1-.2
100=125*.8
100=100